OpenStudy (yinshy):

A cylinder has a base with a radius of 3 cm and a height of 12 cm. What is the volume of the cylinder? (Use π = 3.14.) A. 113.04 cm3 B. 452.16 cm3 C. 339.12 cm3 D. 1356.48cm3

OpenStudy (yinshy):

@KendrickLamar2014

OpenStudy (kendricklamar2014):

Formula for the volume of a cylinder: \[V = \pi \times r^2 \times h\]